The half life of a particular radioactive isotape of thorium is 8 minute. Find an exponential decay equation for this isotope of the form y= ae^-kt where t is in minutes.
Answer by stanbon(75887)

The half life of a particular radioactive isotape of thorium is 8 minute. Find an exponential decay equation for this isotope of the form y= ae^-kt where t is in minutes.
------
y = a*e^(-kt)
---
If "a" is the original amt. of isotope, a/2 is 1/2*a.
---
Solve for "k":
a/2 = a*e^(-kt)
e^(-kt) = 1/2
(-kt) = ln(1/2)
---
(-8k) = -0.6931
k = 0.0866
----
Equation:
y = a*e^(-0.0866t)
=======================
Cheers,
Stan H.